# Sequential Change Detection under Markov Setup with Unknown Prechange and Postchange Distributions

### v2

@@ III-C Behaviour Under Prechange Distribution
$\sum_{b\in\mathcal{X}}\frac{N(b\mid x_{n_{0}+1})}{n}\cdot\log\frac{\mu_{0}(b)}{\widehat{\mu}_{0}(b)}$
@@
$\frac{\mu_{0}(a\mid b)}{{\widehat{\ \mu}}_{0}(a\mid b)}<\frac{\mu_{0}(a\mid b)}{\mu_{0}(a\mid b)-\delta},\frac{\mu_{0}(b)}{{\widehat{\ \mu}}_{0}(b)}<\frac{\mu_{0}(b)}{\mu_{0}(b)-\delta^{\prime}}$
@@
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\frac{n-1}{n}\cdot\log\left(\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta}\right)+\frac{1}{n}\cdot\log(\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}})$
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\frac{(n-1)\log(\beta_{1})+\log(\beta_{2})}{n}=W_{n}$
Lemma 3:The probability of error (false alarm rate) for the proposed test is bounded above by,
@@
where $\epsilon_{0}\to 0$$\beta_{1}\to 1$ and $\beta_{2}\to 1$ as $n_{0}\to\infty$.
@@
$\mu_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)=\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,nf_{n}(x_{n_{0}+1}^{n_{0}+n})}$
@@
$\mu_{0}\!\left(N(\alpha)<\infty\right)=\mu_{0}(B)=\sum_{n=1}^{\infty}\mu_{0}\!\left(B_{n}\cap E_{n}\right)+\mu_{0}\!\left(B\cap(E)^{c}\right)$
$\sum_{n=1}^{\infty}\left(\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}\mu_{0}\!\left(x_{1}^{n_{0}+n}\right)\right)+\mu_{0}\!\left(B\cap(E)^{c}\right)$
As $B_{n}$ consists of only $n$-long sequences, we have $B_{n}\cap E=B_{n}\cap E_{n}$.For first term of (3.18) we have,
@@
$\leq\alpha\cdot 2^{\,-n(\lambda-W_{n})}$
(3.20) follows from (3.16) while the third inequality is due to Lemma 2 as $x_{1}^{n_{0}+n}\in E_{n}$.The fourth inequality comes from definition of $B_{n}$ while the last inequality comes from Kraft’s inequality. As $P(A\cap B)\leq{P(B)}$,and from (3.5) and (3.6), we have,
$\mu_{0}(B\cap(E)^{c})\leq{\mu_{0}((E)^{c})}\leq{\epsilon}+{\epsilon^{\prime}}={\epsilon_{0}}$
Combining both equations and from (3.14) we have,
@@
By comparing the probability of error expressions in Lemma 3 and for actual pre-change case (2.19), we observe that the probability of error increases in the case of the empirical pre-change distribution test. The probability of error now becomes a function of $\log\beta_{1}$$\beta_{1}$ and $\beta_{2}$. For the probability of error to be bounded, we require:
$\lambda>\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))$
In contrast, for the actual pre-change distribution case, $\lambda>0$ would have sufficed. However, as $n_{0}\to\infty$, the lower bound $\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))$,the error term $\epsilon_{0}$ and $\beta_{1},\beta_{2}$ can be made as small as needed.
@@
$\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))<\lambda<D(\mu_{1}||\hat{\mu}_{0})$
III-D Behaviour Under Postchange Distribution
Lemma 4: Let the conditions of Lemma 1 hold true.The proposed test (3.7) will terminate with probability 1 under $\mu_{1}$ if $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$ i.e.,
$\mu_{1}(N(\alpha)<\infty)=1$
Proof:The proof of this lemma is the same as Lemma 5.4 of [25], replacing $\mu_{0}$ with ${\widehat{\mu}}_{0}$.The replacement is valid in this case as the behaviour of the test under $\mu_{1}$ is identical to that of a test where prechange distribution is $\mu_{0}$.
Theorem 2: Let $\mu_{0}$ and $\mu_{1}$ be stationary and ergodic finite order Markov sources satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$. For the Modified CUSUM Test (3.7) with codes satisfying (2.7) and stopping time $N(\alpha)$ (3.10) with $\alpha\to 0$, for $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$, we have
${E_{1}}(N(\alpha)\,|\,x_{1}^{n_{0}})\sim\frac{|\log\alpha|}{D(\mu_{1}||{\widehat{\mu}}_{0})-\lambda}$
Proof:This theorem can be proved by arguing along the same lines as Theorem 5.7 of [25]
@@ IV Performance Bounds
To study the performance of the proposed modified test, we use the following parameter to denote the stopping time of the test,
$M(\gamma)=\inf\left\{n:\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma\right\}$
@@
Consider the case where first m1m-1 samples are generated by $\mu_{0}$ and the rest of the samples come from $\mu_{1}$. In this case, the change point is $m$.Let $P_{m}$} and $E_{m}$} be the corresponding probability distribution and expectation for this case.Then,
$P_{m}\big(x_{n_{0}+1}^{\,n_{0}+n}\big)=\mu_{0}\big(x_{n_{0}+1}^{\,n_{0}+m}\big)\,\mu_{1}\big(x_{n_{0}+m}^{\,n_{0}+n}\big),$
due to the mutual independence of sources before and after the change [16, 25].
For a given stopping time N,Lorden introduced the following quantity (4.2) to serve as minimax criterion for quickness of reaction to a change [6].It states that,
$\overline{E_{1}}(S)=\sup_{m\geq 1}\,\text{ess sup}\,{E_{m}}\left[(S-m+1)^{+}\left|X_{1},\dots,X_{m-1}\right|\right]$
$\overline{E_{1}}(S)$ is the smallest bound on the average no.of differently distributed samples observed before stopping, guaranteed irrespective of behaviour of samples before the change. To study the behaviour of $\overline{E}_{1}(M(\gamma)\,|\,x_{1}^{n_{0}})$, first have to investigate the characteristics of ${E_{0}}(M(\gamma))$, which is the expected time after which the algorithm detects a change and stops the test for the no source change case. Ideally, this has to be $\infty$ as the algorithm should not detect a change when there isn’t any.
Theorem 3 [16]: Lorden’s Theorem was extended for stationary and ergodic processes $(\mu_{0},\mu_{1})$ satisfying Berk’s mixing conditions (2.16),(2.17) such that,
${\mu_{0}}(N<\infty)\leq\alpha$
And for $k=1,2,\dots$, let $N_{k}$ be the stopping variable obtained by applying $N$ to $X_{k},X_{k+1},\dots$, such that,
$N^{\ast}=\min\{N_{k}+k-1\mid k=1,2,\dots\}$
@@
$\overline{E_{1}}(N^{\ast})\leq{E_{1}}(N)$
Define $S_{\gamma}$ to be the class of all sequential change detection tests’ stopping times $S$ satisfying (for $\gamma>0$):
${E_{0}}(S)\geq\frac{1+\gamma}{2}>\frac{\gamma}{2}$
Theorem 4 [16]: For two stationary ergodic and mutually independent stochastic processes with memory $\mu_{0},\mu_{1}$ satisfying (2.10) and Berk’s mixing conditions(2.16),(2.17), as $\gamma\to\infty$, we have,
$\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\sim\frac{\log\gamma}{D(\mu_{1}||\mu_{0})}$
We consider stationary ergodic and mutually independent finite order markov sources $\mu_{0},\mu_{1}$ satisfying (2.10),(2.16),(2.17)
@@
${E_{0}}\left(M(\gamma)\right)\geq\frac{1}{2}+\frac{\gamma}{\frac{2\beta_{2}}{\beta_{1}(2^{\lambda-\log\beta_{1}}-1)}+2\epsilon_{0}\gamma}>\frac{\gamma}{\frac{2\beta_{2}}{\beta_{1}(2^{\lambda-\log\beta_{1}}-1)}+2\epsilon_{0}\gamma}$
@@
$\overline{E_{1}}(M(\gamma)\mid x_{1}^{n_{0}})\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}$
Proof: Equation (4.9) follows from (3.15) of Lemma 3 and (4.5) of Theorem 3(restatement of Theorem 1 from [16]). Equation (4.10) follows from (4.6) of Theorem 3(restatement of Theorem 1 from [16]) and Theorem 2.
Theorem 6:Let $\mu_{0}$ and $\mu_{1}$ be stationary and ergodic finite order Markov sources satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$.For the Modified CUSUM Test (3.7), we can choose $\lambda$ and $\eta$ such that stopping time $M(\eta)=M\in S_{\gamma}$ for universal codes satisfying (2.7) and as $\gamma\to\infty$ we have,
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\inf_{s\in S_{\gamma}}\overline{E_{1}}(S)\quad\text{a.s.}\ [\mu_{0}]$
@@
Consider $M=M(\eta)$, Then by Lemma 3 and Theorem 3 (or equivalently by Theorem 5), we have,
@@
[…]
${\lambda}=D(\mu_{1}\|\hat{\mu}_{0})-\frac{D(\mu_{1}\|\mu_{0})}{1+\vartheta}$
@@
$\overline{E_{1}}(M\mid x_{1}^{n_{0}})\sim\frac{\log\eta}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}$
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}\quad\text{a.s.}\ [\mu_{0}]$
@@
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\frac{\log\gamma}{D(\mu_{1}||{\mu_{0}})}\quad\text{a.s.}\ [\mu_{0}]$
@@
$\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\sim\frac{\log\gamma}{D(\mu_{1}||\mu_{0})}$
Hence, as $n_{0}\to\infty$ and because samples $x_{1}^{n_{0}}$ are generated using $\mu_{0}$, we get
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\inf_{s\in S_{\gamma}}\overline{E_{1}}(S)\quad\text{a.s.}\ [\mu_{0}]$
@@ V Conclusion and Future Work
In this paper, we have extended the results of 2022 for stationary and ergodic finite order Markov sources under suitable conditions.One can aim to generalise these results further under Lai’s framework [20] who has weakened the requirement of independence of observations before and after change to obtain an extension of Lorden’s result.The performance of proposed test can also be studied under other alternate criteria introduced by Lai [20].We have evaluated under Lorden’s criteria as our work extends [24, 25, 26] all of which are studied under Lorden’s criteria.A natural direction for future work would be to study under Pollack’s criteria[14] where variants for unknown distributions exist and compare performance.

### v3

@@ III-C Behaviour Under Pre-change Distribution
$\sum_{b\in\mathcal{X}}\frac{N(b\mid x_{n_{0}+1})}{n}\cdot\log\frac{\mu_{0}(b)}{\widehat{\mu}_{0}(b)}.$
@@
$\frac{\mu_{0}(a\mid b)}{{\widehat{\ \mu}}_{0}(a\mid b)}<\frac{\mu_{0}(a\mid b)}{\mu_{0}(a\mid b)-\delta},\frac{\mu_{0}(b)}{{\widehat{\ \mu}}_{0}(b)}<\frac{\mu_{0}(b)}{\mu_{0}(b)-\delta^{\prime}}.$
@@
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\frac{n-1}{n}\cdot\log\left(\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta}\right)+$
$\frac{1}{n}\cdot\log(\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}}).$
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\frac{(n-1)\log(\beta_{1})+\log(\beta_{2})}{n}=W_{n}.$
Lemma 3: The probability of error for the proposed test is bounded above by,
@@
where $\epsilon_{0}\to 0$, $\beta_{1}\to 1$ and $\beta_{2}\to 1$ as $n_{0}\to\infty$.
@@
$\mu_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)=\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,nf_{n}(x_{n_{0}+1}^{n_{0}+n})}.$
@@
$\mu_{0}\!\left(N(\alpha)<\infty\right)=\mu_{0}(B)=\sum_{n=1}^{\infty}\mu_{0}\!\left(B_{n}\cap E_{n}\right)+$
$\mu_{0}\!\left(B\cap(E)^{c}\right).$
$\sum_{n=1}^{\infty}\left(\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}\mu_{0}\!\left(x_{1}^{n_{0}+n}\right)\right)+\mu_{0}\!\left(B\cap(E)^{c}\right).$
As $B_{n}$ consists of only $n$-long sequences, we have $B_{n}\cap E=B_{n}\cap E_{n}$. For first term of (38) we have,
@@
$\leq\alpha\cdot 2^{\,-n(\lambda-W_{n})}.$
(40) follows from (36) while the third inequality is due to Lemma 2 as $x_{1}^{n_{0}+n}\in E_{n}$. The fourth inequality comes from definition of $B_{n}$ while the last inequality comes from Kraft’s inequality. As $P(A\cap B)\leq{P(B)}$, and from (25) and (26), we have,
$\mu_{0}(B\cap(E)^{c})\leq{\mu_{0}((E)^{c})}\leq{\epsilon}+{\epsilon^{\prime}}={\epsilon_{0}}.$
Combining both equations and from (34) we have,
@@
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{\dfrac{\beta_{1}\left(2^{\lambda-\log\beta_{1}}-1\right)}{\beta_{2}}}+\epsilon_{0}.$
By comparing the probability of error expressions in Lemma 3 and for known pre-change case (18), we observe that the probability of error increases for our test. The probability of error now becomes a function of $\log\beta_{1}$, $\beta_{1}$ and $\beta_{2}$. For the probability of error to be bounded, we require:
$\lambda>\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0})).$
In contrast, for known pre-change case (18), $\lambda>0$ would have sufficed. However, as $n_{0}\to\infty$, the lower bound $\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))$, the error term $\epsilon_{0}$ and $\beta_{1},\beta_{2}$ can be made as small as needed.
@@
$\max(\log\beta_{1},D(\mu_{0}||{\widehat{\mu}}_{0}))<\lambda<D(\mu_{1}||\hat{\mu}_{0}).$
III-D Behaviour Under Post-change Distribution
Lemma 4: Let the conditions of Lemma 1 hold true. The proposed test (27) will terminate with probability 1 under $\mu_{1}$ if $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$, i.e.,
$\mu_{1}(N(\alpha)<\infty)=1.$
Proof: The proof of this lemma is the same as that of Lemma 3 in [24], since its proof is dependent on Lemma 1, replacing $\mu_{0}$ with ${\widehat{\mu}}_{0}$. The replacement is valid in this case as the behaviour of the test under $\mu_{1}$ is identical to that of a test where pre-change distribution is $\mu_{0}$.
Theorem 2: Let $\mu_{0}$ and $\mu_{1}$ be stationary, ergodic first-order Markov sources satisfying (9), (15), (16) and $\mu_{1}\ll\mu_{0}$. For the Modified CUSUM Test (27) with a universal code satisfying (5) and stopping time $N(\alpha)$ (30) as $\alpha\to 0$, for $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$, we have,
${E_{1}}(N(\alpha)\,|\,x_{1}^{n_{0}})\sim\frac{|\log\alpha|}{D(\mu_{1}||{\widehat{\mu}}_{0})-\lambda}.$
Proof: This theorem can be proved by arguing along similar lines as Theorem 2 of [24] since its proof is dependent on (3), (9) which are satisfied by the codes and sources considered in this paper.
@@ IV Performance Bounds
To study the performance of the proposed modified test, we use $M(\gamma)$ to denote the stopping time of the test,
$M(\gamma)=\inf\Bigl\{n:\max_{1\leq k\leq n}\Bigl($ $-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+k}^{n_{0}+n})$ (52)
$-n\lambda\Bigr)\geq\log\gamma\Bigr\}.$
@@
Consider the case where first m1m-1 samples are generated by $\mu_{0}$ and the rest of the samples come from $\mu_{1}$. In this case, the change point is $m$. Let $P_{m}$} and $E_{m}$} be the corresponding probability distribution and expectation for this case. Then, $P_{m}\big(x_{1}^{\,n_{0}+n}\big)=\mu_{0}\big(x_{1}^{\,m-1}\big)\,\mu_{1}\big(x_{m}^{\,n_{0}+n}\big)$ due to the mutual independence of sources before and after the change [16, 25].
For a given stopping time $S$, Lorden proposed the following quantity (53) to serve as minimax type of criterion for quickness of reaction to a change [6].
$\overline{E_{1}}(S)=\sup_{m\geq 1}\,\text{ess sup}\,{E_{m}}\left[(S-m+1)^{+}\left|X_{1},\dots,X_{m-1}\right|\right].$
To investigate the behaviour of $\overline{E}_{1}(M(\gamma)\,|\,x_{1}^{n_{0}})$, we first study the characteristics of ${E_{0}}(M(\gamma))$, which is the expected time after which the algorithm detects a change and stops the test for the no source change case. Ideally, this has to be $\infty$ as the algorithm should not detect a change when there isn’t any.
Theorem 3 [16]: Lorden’s Theorem was extended for stationary and ergodic processes $(\mu_{0},\mu_{1})$ satisfying Berk’s mixing conditions (15), (16) such that,
${\mu_{0}}(N<\infty)\leq\alpha.$
For $k=1,2,\dots$, let $N_{k}$ be the stopping variable obtained by applying $N$ to $X_{k},X_{k+1},\dots$, such that,
$N^{\ast}=\min\{N_{k}+k-1\mid k=1,2,\dots\}.$
@@
$\overline{E_{1}}(N^{\ast})\leq{E_{1}}(N).$
Define $S_{\gamma}$ to be the class of all sequential change detection tests’ stopping times $S$ satisfying:
${E_{0}}(S)\geq\frac{1+\gamma}{2}>\frac{\gamma}{2}.$
Theorem 4 [16]: For two stationary, ergodic and mutually independent stochastic processes with memory $\mu_{0},\mu_{1}$ satisfying Berk’s mixing conditions (15), (16), as $\gamma\to\infty$, we have,
$\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\sim\frac{\log\gamma}{D(\mu_{1}||\mu_{0})}.$
@@
${E_{0}}\left(M(\gamma)\right)\geq\frac{1}{2}+\frac{\gamma}{\frac{2\beta_{2}}{\beta_{1}(2^{\lambda-\log\beta_{1}}-1)}+2\epsilon_{0}\gamma}$ (60)
$>\frac{\gamma}{\frac{2\beta_{2}}{\beta_{1}(2^{\lambda-\log\beta_{1}}-1)}+2\epsilon_{0}\gamma},$
@@
$\overline{E_{1}}(M(\gamma)\mid x_{1}^{n_{0}})\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}.$
Proof: Equation (60) follows from (35) of Lemma 3 and (56) of Theorem 3 (restatement of Theorem 1 from [16]). Equation (61) follows from (57) of Theorem 3 (restatement of Theorem 1 from [16]) and Theorem 2.
Theorem 6: Let $\mu_{0}$ and $\mu_{1}$ be stationary, ergodic and mutually independent first-order Markov sources satisfying (9), (15), (16) and $\mu_{1}\ll\mu_{0}$. For the Modified CUSUM Test (27) with universal code satisfying (5), we can choose $\lambda$ and $\eta$ such that stopping time $M(\eta)=M\in S_{\gamma}$ and as $\gamma\to\infty$ we have,
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\quad\text{a.s.}\ [\mu_{0}],$
@@
Consider $M=M(\eta)$. By Theorem 5 (or equivalently by Lemma 3 and Theorem 3), we have,
@@
Hence, $M\in S_{\gamma}$. For a choice of $\vartheta>0$ let $\lambda$ be chosen such that,
${\lambda}=D(\mu_{1}\|\hat{\mu}_{0})-\frac{D(\mu_{1}\|\mu_{0})}{1+\vartheta}.$
@@
$\overline{E_{1}}(M\mid x_{1}^{n_{0}})\sim\frac{\log\eta}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}.$
Hence,
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}\quad\text{a.s.}\ [\mu_{0}].$
@@
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\frac{\log\gamma}{D(\mu_{1}||{\mu_{0}})}\quad\text{a.s.}\ [\mu_{0}].$
@@
$\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\sim\frac{\log\gamma}{D(\mu_{1}||\mu_{0})}.$
Hence, as $n_{0}\to\infty$ and because samples $x_{1}^{n_{0}}$ are generated using $\mu_{0}$, we get,
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\quad\text{a.s.}\ [\mu_{0}].$
@@ V Conclusion and Future Work
The analysis holds for any stationary, ergodic, mutually independent finite-order Markov sources (pre-change distribution estimation, Lemma 2, Lemma 3 and equation (60) can be modified as per Markov source order). Hence, we have extended the results of 2022 for stationary and ergodic finite-order Markov sources under suitable conditions. One can aim to generalise these results further under Lai’s framework [20] who has weakened the requirement of independence of observations before and after change to obtain an extension of Lorden’s result. The performance of proposed test can also be studied under other alternate criteria introduced by Lai [20]. We have evaluated under Lorden’s criteria as our work extends [24, 25, 26] all of which are studied under Lorden’s criteria.